a) A direct current generator of 127V provides 2.9kW to an electric oven. Determine the current it provides. Find the resistance. 

b) A wire 2 mm diameter and 100 m length has a resistance of 3Ω (at 20°C). Find its resistivity.



1
Expert's answer
2021-11-01T19:24:51-0400

a)

"W=VI \\

I = \frac{W}{V} = \frac{2900}{127} = 22.83 \;mp

V=IR \\

R = \frac{V}{I} = \frac{127}{22.83} = 5.56 \; \Omega"

b) L = 100 m

R = 3 "\Omega"

d = 2 mm = 0.002 m

r = 1 mm = 0.001 m

"A = \pi r^2 \\

A = 3.14 \times 10^{-6} \;m^2 \\

p = \frac{RA}{L} \\

p = \frac{3 \times 3.14 \times 10^{-6}}{100} = 9.42 \times 10^{-8} \; \Omega m"